OpenStudy (anonymous):

what does y=f^-1(x) mean?

OpenStudy (anonymous):

I think it means the inverse function. Does that make sense?

OpenStudy (anonymous):

Is it y=f^(-1)x or y =f^(-1x)

OpenStudy (anonymous):

so a reflection right?

OpenStudy (anonymous):

nope its the inverse. Example y=2x the inverse would be y=1/2x

OpenStudy (anonymous):

the -1 is outside the bracket and its an exponent

OpenStudy (anonymous):

Inverse function notation looks like\[f ^{-1}(x)\]